# Divide-and-Conquer Meets Consensus: Unleashing the Power of Functions in Code Generation

### v1

@@ 1 Introduction
We conduct extensive experiments on code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Khan et al., 2023) with GPT (Ouyang et al., 2022; OpenAI, 2023), outperforming state-of-the-art methods by $\mathbf{+9.8\%}$ on average. Experiments are further carried out on the mathematical competition benchmark, MATH (Hendrycks et al., 2021b), achieving a $\mathbf{+6.0}$ improvement with GPT-4, indicating that FunCoder can also generalize to complex reasoning. Our method is observed to be equally effective on open-source models (Rozière et al., 2023; Pinnaparaju et al., 2024; Meta AI, 2024), with an average gain over baseline of $\mathbf{+38.0\%}$ on HumanEval and $\mathbf{+61.1\%}$ on MATH. Additional analysis also shows the advantage of both divide-and-conquer and functional consensus.
@@ 2 FunCoder: Divide-and-Conquer Meets Consensus
Figure 2: Left: Algorithm for FunCoder procedure. Right: Comparison between decomposition by planning and our approach. FunCoder introduces new functions to describe sub-goals solely with code, achieving a more natural way of requirement decomposition.
2.1 Divide-and-Conquer for Iterative Programming
@@ 2.2 Functionality Similarity as a Consensus
The decomposition of complex tasks benefits from solving easier sub-goals, but might introduce the risks of cascading errors. To mitigate this, we introduce Functional Consensus which aims at reducing inconsistencies in program behavior. This is achieved by sampling multiple functions and selecting the one that exhibits consensus, as measured by the aggregated similarity of functionality between candidates, thus abating outlier functionalities.
Functionality Similarity  A program specifies its functionality (or behavior) through the control flow defined by its code semantics. However, comparing the functionalities between two programs based on their semantics is somewhat challenging. By decomposing the requirement into functions, FunCoder is able to view the function behavior as a black box that maps arguments into return values. Considering two functions $f$ and $g$ with the same input domain $D(f)=D(g)$, we define the similarity between them $sim(f,g)$ as the identicalness of outputs when given the same input values.
@@ 3.1.1 Experiment Setup
GPT-3.5 Standard 68.3 - 72.0 - 44.4 15.2 4.6 0.0 20.2
CodeT 81.1 +12.8 76.0 +4.0 50.6 16.1 8.0 0.0 23.2
Reflexion 69.5 +1.2 72.5 +0.5 44.4 17.0 5.7 0.0 20.6
LDB 82.9 +14.6 76.0 +4.0 - - - - -
FunCoder 85.4 +17.1 78.5 +6.5 62.4 29.5 11.6 0.0 31.4
@@ 3.1.2 Results
Table 2: Code generation performance with open-source models on HumanEval.
Model Category Param Standard CodeT Reflexion FunCoder
Llama3 Text/Chat 8B 61.6 68.9 59.1 79.9 (+11.0)
StableCode Code 3B 61.0 75.0 61.6 81.0 (+6.0)
CodeLlama Code 34B 43.9 55.5 41.5 66.5 (+11.0)
Evaluation is also performed over community LLMs, Llama3 (Meta AI, 2024), StableCode (Pinnaparaju et al., 2024), and CodeLlama (Rozière et al., 2023) with results in Table 210. FunCoder consistently boosts the performance of smaller models in code generation, with an averaged improvement of +38.0% compared to standard prompting, and outperforms the previous best method CodeT by +14.6% on HumanEval. Experiment results demonstrate that our method archives state-of-the-art performance on various models, ranging from basic programming to competition contests.
@@ 3.1.3 Analysis
FunCoder Democratize to Smaller LLMs  Limited by the LLM capabilities, the application of self-improvement or multi-agent methods on smaller models is without ease. By keeping decomposition and composition within the code generation process, our approach exhibits better generalization. As shown in Table 12, with FunCoder, Llama38b and StableCode3b achieve around 1.18×1.18$\times$1.18 × performance to standard GPT-3.5, and are closely aligned with GPT-4 by about $97\%$ on HumanEval.
@@
FunCoder@5 Consensus@5 83.5 (+15.2) 4040.8
@@
Ablation and Token Usage  To analyze the impact of dividing, conquering, and functional consensus in FunCoder, we carry out an ablation study with different settings. A study that replaces consensus with self-testing is also included. The ablation is constructed on HumanEval with GPT-3.5, as shown in Table 3. We observe that function decomposition and re-composition deliver cumulative performance improvements. Functional consistency is also shown to prevail over self-testing. Putting them all together, FunCoder received a $+17.1$ improvement with 5.09×5.09$\times$5.09 × more tokens over baseline. Compared to previous SOTA LDB ($\approx 23$K tokens), we are able to gain $+2.5$ in performance with $76.5\%$ token usage reduction.
@@ 4 Related Work
Large Language Model for Code  Code pre-training has received widespread attention, with early models based on small language models (SLM) (Feng et al., 2020; Lu et al., 2021; Wang et al., 2021). In recent years, with the development of large-scale pre-training techniques, code LLM has emerged, showing remarkable performance in downstream code tasks (Chen et al., 2021; Nijkamp et al., 2023; Li et al., 2022; Rozière et al., 2023; Li et al., 2023b; Guo et al., 2024). Tasks between code and natural language (NL) can be generally divided into three major categories: NL2Code tasks such as code generation (Austin et al., 2021; Chen et al., 2021; Hendrycks et al., 2021a; Khan et al., 2023) and code search (Husain et al., 2019a); Code2Code tasks including code completion (Lu et al., 2021; Zhang et al., 2023a; Liu et al., 2024), code translation (Ahmad et al., 2023; Zhu et al., 2022; Yan et al., 2023), and test generation (Siddiq et al., 2023; Schäfer et al., 2024); Code2NL tasks like code summarization (Husain et al., 2019b; Jin et al., 2023). This paper focuses on code generation tasks, ranging from basic to competition level.
Code Refinement and Self-Testing  Code doesn’t always run as expected; it could contain syntax errors, dead loops, or bugs. It’s essential to debug and refine the code to ensure better quality. CodeT (Chen et al., 2023a) generates unit-tests to score the implementation. Self-improvement methods (Madaan et al., 2023; Shinn et al., 2023; Chen et al., 2024; Zhong et al., 2024) design closed-loop procedures that repeatedly refine the code based on the feedback. Like real-life software development processes, multi-agent frameworks (Hong et al., 2024; Qian et al., 2023) construct specific LLM roles, Tester or QA to generate tests. These studies adopt a shared paradigm wherein self-tests are generated through LLMs. However, Olausson et al. (2024) points out the challenge that LLMs have certain shortcomings in self-repairing their code. This paper avoids these shortcomings by proposing functional consensus as a reliable method of evaluation.
@@ A.1 Implementation Details
We access the OpenAI models GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-1106-preview) through the Azure API. Weights of community models Llama3 (Meta-Llama-3-8B-Instruct), StableCode (stable-code-instruct-3b), and CodeLlama (CodeLlama-34b-Instruct-hf) are downloaded from HuggingFace (Wolf et al., 2019) and are served over an OpenAI-like API on a single A100-80G GPU under BF16 precision with vLLM (Kwon et al., 2023).
@@ A.5 Details of Analysis
(Figure 3.a) The preliminary study is conducted on the HumanEval dataset, which includes system tests $S$ to evaluate the accuracy of the program, as well as one human-annotated canonical solution $c$. For each question, we: (1) Obtain one solution program $p$ from Standard Prompting. (2) Prompt the model to generate 7 self-tests $T$ based on the question and entry function. The self-test is in the form of the unit test assert f(x) == y. We then judge the generated program $p$ and canonical solution $c$ over the self-tests $T$ and system tests $S$. Formally, a pair $(x,Y)$ is used to identify whether program $x$ passes test $Y$. Where $(p,S)$ indicates that the program can pass the system tests, demonstrating its correctness. And $\neg(c,T)$ means the canonical solution can not pass self-tests, suggesting that the tests generated by model could be wrong. The self-test results on generated programs are first evaluated and divided into two classes: self-test passed or failed. If the self-test passes, the self-improvement methods will stop iteration and pick this program as a final result. The next step is to determine whether the program can pass system tests. If the self-test fails, it indicated that there could be an error in the program or test itself. In this case, the correctness of the program is checked using final tests $(p,S)$ and the correctness of the unit test by canonical program $(c,T)$. The results on GPT-3.5 and StableCode are shown in Figure 3 and detailed explanations about these conditions can be found in Table 9.
@@
A.6 Supplementary Results
Open-source model on MBPP and xCodeEval
We provide the results for community models Llama3, StableCode, and CodeLlama on other code generation benchmarks in Table 10. Our method consistently improves the baseline by averaging $12\%$ on MBPP and $197\%$ on xCodeEval. It is worth noting that the small model has a low pass rate on the competition problems, leading to a relatively high randomness, so we report the median by running 3 experiments.
Table 10: Results for MBPP and xCodeEval with community models.
Model Method MBPP xCodeEval
Pass@1 $\Delta\uparrow$ Easy Mid Hard Expert All
Llama38b Standard 60.5 - 9.0 1.8 0.0 0.0 3.6
FunCoder 62.5 +2.0 22.0 0.9 0.0 0.0 8.0
StableCode3b Standard 51.5 - 7.3 0.9 0.0 0.0 2.8
FunCoder 63.5 +12.0 13.5 4.5 1.1 0.0 6.2
CodeLlama34b Standard 53.5 - 2.3 0.0 0.0 0.0 0.8
FunCoder 58.5 +5.0 10.2 0.0 0.0 0.0 3.6
Token Usage on Other Methods

### v2

@@ 1 Introduction
We conduct extensive experiments on code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Khan et al., 2023) with GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), outperforming state-of-the-art methods by $\mathbf{+9.8\%}$ on average. Experiments are further carried out on the mathematical competition benchmark, MATH (Hendrycks et al., 2021b), achieving a $\mathbf{+6.0}$ improvement with GPT-4, indicating that FunCoder can also generalize to complex reasoning. Our method is observed to be equally effective on open-source models (Meta AI, 2024; Mistral AI, 2024; Pinnaparaju et al., 2024; Rozière et al., 2023; Lozhkov et al., 2024), with an average gain over baseline of $\mathbf{+31.5\%}$ on HumanEval and $\mathbf{+47.7\%}$ on MATH. Additional analysis also shows the advantage of both divide-and-conquer and functional consensus. Our code is made openly available at https://github.com/cometeme/funcoder.
@@ 2.1 Divide-and-Conquer for Iterative Programming
Figure 2: Left: Algorithm for FunCoder, explained in detail in Appendix A.6. Right: Comparison between decomposition by planning and our approach. FunCoder introduces new functions to describe sub-goals solely with code, achieving a more natural way of requirement decomposition.
@@ 2.2 Functionality Similarity as a Consensus
The decomposition of complex tasks benefits from solving easier sub-goals, but might introduce the risks of cascading errors, which refers to errors in sub-functions that lead to errors in ancestral functions. To mitigate this, we introduce Functional Consensus which aims at reducing inconsistencies in program behavior. This is achieved by sampling multiple functions and selecting the one that exhibits consensus, as measured by the aggregated similarity of functionality between candidates, thus abating outlier functionalities.
Functionality Similarity  A program specifies its functionality (or behavior) through the control flow and logic defined by its code semantics. However, comparing the functionalities between two programs based on their semantics is somewhat challenging. By decomposing the requirement into functions, FunCoder is able to view the function behavior as a black box that maps arguments into return values. Considering two functions $f$ and $g$ with the same input domain $D(f)=D(g)$, we define the similarity between them $sim(f,g)$ as the identicalness of outputs when given the same input values.
@@ 3.1.1 Experiment Setup
Llama38b Standard 61.6 - 60.5 - 9.0 1.8 0.0 0.0 3.6
CodeT 68.9 +7.3 61.5 +1.0 12.4 0.0 0.0 0.0 4.4
FunCoder 79.7 +18.1 62.5 +2.0 22.0 0.9 0.0 0.0 8.0
StableCode3b Standard 61.0 - 51.5 - 7.3 0.9 0.0 0.0 2.8
CodeT 75.0 +14.0 57.5 +6.0 11.2 1.8 0.0 0.0 4.6
FunCoder 81.0 +20.0 63.5 +12.0 13.5 4.5 1.1 0.0 6.2
CodeLlama34b Standard 43.9 - 53.5 - 2.3 0.0 0.0 0.0 0.8
CodeT 55.5 +11.6 56.5 +3.0 10.1 0.0 0.0 0.0 3.6
FunCoder 66.5 +22.6 58.5 +5.0 10.2 0.0 0.0 0.0 3.6
@@ 3.1.2 Results
Evaluation is also performed over community LLMs, Llama3 (Meta AI, 2024), StableCode (Pinnaparaju et al., 2024) and CodeLlama (Rozière et al., 2023) with results in Table 1. FunCoder consistently boosts the performance of smaller models in code generation, demonstrating notable improvements compared to standard prompting on HumanEval, which gained +29.4% on Llama3, +32.8% on StableCode, and even +51.5% on CodeLlama, outperforming that from the previous best method CodeT. We also supplement results on GPT-4o mini, Codestral and StarCoder2 in Table 11. Experiment results demonstrate that our method archives state-of-the-art performance on various models, ranging from basic programming to competition contests.
@@ 3.1.3 Analysis
FunCoder Democratize to Smaller LLMs  Limited by the LLM capabilities, the application of self-improvement or multi-agent methods on smaller models is without ease. By keeping decomposition and composition within the code generation process, our approach exhibits better generalization. As shown in Table 1, with FunCoder, StableCode3b achieves around $118.6\%$ relative performance to standard GPT-3.5, and also aligns closely with GPT-4 by about $97.7\%$ on HumanEval.
@@
Two-pass + CL@11 Clustering@11 75.0 (+6.7) 5070.7
FunCoder@5 Consensus@5 83.5 (+15.2) 4040.9
@@
Ablation and Token Usage  To analyze the impact of dividing, conquering, and functional consensus in FunCoder, we carry out an ablation study with different settings. Studies that replace consensus with self-testing, or with AlphaCode-like (Li et al., 2022) clustering, are also included. The ablation is constructed on HumanEval with GPT-3.5, as shown in Table 2. Note that to generate every program FunCoder costs only $O(kN)$ tokens, where $k$ is the number of sampled candidates, and $N$ is the token length of the final program. This is further exemplified and explained in §A.7. We observe that function decomposition and re-composition deliver cumulative performance improvements. Functional consistency is also shown to prevail over self-testing. Putting them all together, FunCoder received a $+17.1$ improvement with just 5.09×5.09$\times$5.09 × more tokens over baseline. Compared to previous SOTA LDB ($\approx 23$K tokens), we are able to gain $+2.5$ in performance with $76.5\%$ token usage reduction.
3.2 Mathematical Reasoning
3.2.1 Experiment Setup
@@ 4 Related Work
Large Language Model for Code  Code pre-training has received widespread attention, with early models based on small language models (SLM) (Feng et al., 2020; Lu et al., 2021; Wang et al., 2021). In recent years, with the development of large-scale pre-training techniques, code LLM has emerged, showing remarkable performance in downstream code tasks (Chen et al., 2021; Nijkamp et al., 2023; Li et al., 2022; Rozière et al., 2023; Li et al., 2023b; Guo et al., 2024). Tasks between code and natural language (NL) can be generally divided into three major categories: NL2Code tasks such as code generation (Austin et al., 2021; Chen et al., 2021; Hendrycks et al., 2021a; Khan et al., 2023) and code search (Husain et al., 2019); Code2Code tasks including code completion (Lu et al., 2021; Zhang et al., 2023a; Liu et al., 2024), code translation (Ahmad et al., 2023; Zhu et al., 2022; Yan et al., 2023), and test generation (Siddiq et al., 2023; Schäfer et al., 2024); Code2NL tasks like code summarization (Husain et al., 2019; Jin et al., 2023). This paper focuses on code generation tasks, ranging from basic to competition level.
Code Refinement and Self-Testing  Code doesn’t always run as expected; it could contain syntax errors, dead loops, or bugs. It’s essential to debug and refine the code to ensure better quality. CodeT (Chen et al., 2023a) generates unit-tests to score the implementation. AlphaCode (Li et al., 2022) clusters programs based on whether generated program outputs were identical or not. Self-improvement methods (Madaan et al., 2023; Shinn et al., 2023; Chen et al., 2024; Zhong et al., 2024) design closed-loop procedures that repeatedly refine the code based on the feedback. Like real-life software development processes, multi-agent frameworks (Hong et al., 2024; Qian et al., 2023) construct specific LLM roles, Tester or QA to generate tests. These studies adopt a shared paradigm wherein self-tests are generated through LLMs. However, Olausson et al. (2024) points out the challenge that LLMs have certain shortcomings in self-repairing their code. This paper avoids these shortcomings by proposing functional consensus as a reliable method of evaluation.
@@ 6 Conclusion
Acknowledgments
We would like to acknowledge the reviewers and chairs for their inspiring and constructive feedback. The research in this article is supported by the National Key Research and Development Project (2021YFF0901602), the National Science Foundation of China (U22B2059, 62276083). Bing Qin and Qianglong Chen are the corresponding authors.
@@ Appendix A Appendix
AlphaCode-like Clustering - AlphaCode proposed a clustering process that elects candidate program from a number of samples, recognizing programs that produce exactly identical outputs as equivalent, and picks one program from the largest cluster.
@@ A.1 Implementation Details
We access the OpenAI models GPT-3.5 (gpt-3.5-turbo-0613), GPT-4 (gpt-4-1106-preview) and GPT-4o mini (gpt-4o-mini-2024-07-18) through Azure OpenAI. Weights of community models Llama3 (Meta-Llama-3-8B-Instruct), Codestral (Codestral-22B-v0.1), StableCode (stable-code-instruct-3b), CodeLlama (CodeLlama-34b-Instruct-hf) and StarCoder2 (starcoder2-15b-instruct-v0.1) are downloaded from HuggingFace (Wolf et al., 2019) and served over an OpenAI-like API on a single A100-80G GPU under BF16 precision with vLLM (Kwon et al., 2023).
@@ A.2 Baseline Details
AlphaCode-like Clustering  is introduced with the model AlphaCode (Li et al., 2022), and samples multiple code solutions $X$. A fine-tuned model is used to generate test inputs upon which programs are evaluated. Programs are then clustered by whether the outputs are (exactly) identical, in which the largest group was selected. The final result comes from this largest group, where any result within would suffice. In our ablation study, we use similar settings to FunCoder, sampling 11 candidate solutions with 0.8 temperature and generate sample inputs likewise.
@@ A.5 Details of Analysis
(Figure 3.a) The preliminary study is conducted on the HumanEval dataset, which includes system tests $S$ to evaluate the accuracy of the program, as well as one human-annotated canonical solution $c$. For each question, we: (1) Obtain one solution program $p$ from Standard Prompting. (2) Prompt the model to generate 7 self-tests $T$ based on the question and entry function. The self-test is in the form of the unit test assert f(x) == y. We then judge the generated program $p$ and canonical solution $c$ over the self-tests $T$ and system tests $S$. Formally, a pair $(x,Y)$ is used to identify whether program $x$ passes test $Y$. Where $(p,S)$ indicates that the program can pass the system tests, demonstrating its correctness. And $\neg(c,T)$ means the canonical solution can not pass self-tests, suggesting that the tests generated by model could be wrong. The self-test results on generated programs are first evaluated and divided into two classes: self-test passed or failed. If the self-test passes, the self-improvement methods will stop iteration and pick this program as a final result. The next step is to determine whether the program can pass system tests. If the self-test fails, it indicates that there could be an error in the program or test itself. In this case, the correctness of the program is checked using final tests $(p,S)$ and the correctness of the unit test by canonical program $(c,T)$. The results on GPT-3.5 and StableCode are shown in Figure 3 and detailed explanations about these conditions can be found in Table 8.
@@
A.6 Detailed Explanation of Algorithm
We hereby provide a detailed explanation of FunCoder algorithm works, with respect to Algorithm 1 from Figure 2 (a copy is included below for simple reading). As mentioned, FunCoder is a recursive process following a DFS pattern. We use square brackets (e.g. [L1]) below to denote line numbers in the pseudocode.
2:Large language model, LLM
3:function FunCoder($f_{\textrm{cur}}$)
4:        — Divide —
6:        for $f_{i}\in\{f_{i}\}$ do
7:               if $b_{i}$ is NotImplemented then
8:                       $f^{*}_{i}\leftarrow\textsc{FunCoder}{}(f_{i})$triangleright recursion
9:               end if
10:               $\textsc{AddChild}(f_{\textrm{cur}},f^{*}_{i})$
11:        end for
12:        — Conquer —
15:        return $f^{*}_{\textrm{cur}}$
16:end function
Figure 5: Left: Algorithm for FunCoder. Right: Decomposition example of A[B[DE]C].
FunCoder [L1], when solving each function $f$, first performs the Divide stage [L3-L9], where the LLM initially writes the function and identifies some potential sub-problems, represented as sub-function stubs (e.g., def f(xs: list[int]) -> int) [L3]. In this process, we identify the sub-problems of the current problem, thereby understanding the dependency between functions.
For each decomposed sub-problem $g_{1},g_{2},\ldots$, we recursively use FunCoder to obtain the final implementation $G_{i}$ for that sub-problem [L5-L8]. This $G_{i}$ shall replace the previously incomplete subfunction stub signature in the final program.
Now that all sub-problems of $f$ are implemented, we move on to the Conquer stage [L11-13] to complete the larger problem. By combining the signature $f$ and the final implementations $G_{1},G_{2},\ldots$ of sub-problems, we generate the complete implementation $F$ [L11] and return it [L13].
We hierarchically describe how this algorithm works in detail by combining it with the example given in the right half of Figure 5.
[a.1] FunCoder(a)
|  [a.3] LLM(a) -> A, {b, c}       # divide
|--[b.1] FunCoder(b)
|  |  [b.3] LLM(b) -> B, {d, e}    # divide
|  |--[d.1] FunCoder(d)
|  |  |  [d.3] LLM(d) -> D, {}     # divide
|  |  +--[d.13] return D           # nothing to conquer
|  |--[e.1] FunCoder(e)
|  |  |  [e.3] LLM(e) -> E, {}     # divide
|  |  +--[e.13] return E           # nothing to conquer
|  |  [b.11] LLM(B, {D, E}) -> B*  # conquer
|  +--[b.13] return B*
|--[c.1] FunCoder(c)
|  |  [c.3] LLM(c) -> C, {}        # divide
|  +--[c.13] return C              # nothing to conquer
|  [a.11] LLM(A, {B, C}) -> A*     # conquer
+--[a.13] return A*                # final result
A.7 Token-cost Complexity
Example
We use the example from Figure 5, where the final program consists of 5 functions A[B[D,E],C], and A serves as the entry to the program. Here we respect the aforementioned notations, and further use the lower-case letter $a$ to represent the number of stub tokens, upper-case $A$ to represent the number of result tokens, for the function A, and other functions likewise. Let $N=A+B+C+D+E$ be the token-cost complexity of the final result.
In Standard, the code is only generated once to complete the given stub. We use parentheses to represent the order of LLM calls in a full process.
(1)           a -> A B C D E
   input tokens  = a
   output tokens = A + B + C + D + E
   overall       = O(N)
In each step of the FunCoder/Divide stage, the to-be-implemented function will serve as the context. The function (stub) will be implemented and sub-function stubs are to be declared.
(1)           a -> A b c
(2)           b -> B d e
(3)           d -> D
(5)           e -> E
(8)           c -> C
   input tokens  = a + b + c + d + e
   output tokens = A + b + B + c + C + d + D + e + E < 2N
   overall       = O(N)
In every FunCoder/Conquer, the context shall include the current function’s definition and finalized implementations of sub-functions. The output is the re-implemented current function.
(4)           d -> kD
(6)           e -> kE
(7)       b D E -> kB
(9)           c -> kC
(10)      a B C -> kA
   input tokens  = a + b + B + c + C + d + D + e + E < 2N
   output tokens = kA + kB + kC + kD + kE = kN
   overall       = O(kN)
These stages in all bring FunCoder’s total token consumption to strictly $O(kN)$ for every problem.
Token Complexity of FunCoder is $O(kN)$
Define $N$ as the token length of the final program, which is correlated to the inherent complexity of the problem, and define $k$ as the number of sampled candidates. We first explain in detail that the worst-case token cost of FunCoder is $O(kN)$:
  • The naive Standard method should naturally generate $O(N)$ tokens. Sampling-based baselines like CodeT cost $O(kN)$ tokens.
  • FunCoder goes through the Divide stage and the Conquer stage for each of the functions.
  • Based on the current function, Divide generates an implementation of itself and stubs for sub-functions. Within this stage, each function would appear at most once in input and twice in output. All Divide stages consume no more than $3N$ tokens.
  • Conquer regenerates the parent function based on its stub and all finalized sub-functions. Herein each function will appear at most twice in input, and sampled $k$ times in output. If $k=1$, consensus is implicitly disabled. All Conquer stages shall consume at most $(k+2)N$ tokens.
So FunCoder requires no more than $(k+5)N$ tokens in input-output, making its token consumption $O(kN)$ even at worst-case, aligning with other sampling-based methods such as CodeT and AlphaCode-like clustering. Furthermore, when sampling is disabled ($k=1$), our method has a token consumption of $O(N)$, which also aligns with the vanilla Standard method.
A.8 Discussion About Functional Consensus
This section focuses on why functional consensus might enhance the correctness of programs and how it differs from other consistency-based methods. Self-consistency (Wang et al., 2023) is widely employed in the realm of LLM reasoning. It samples multiple sets of answers and uses voting to select the most consistent result, where the answers typically consist of named entities, choice options, or numbers. However, this approach faces challenges when voting on sampled programs, as programs describe executable logic instead of data, making it unobvious to determine whether two programs are equivalent just from the looks.
When it comes to picking programs, functional consensus in FunCoder looked beyond the literal symbols and used a different approach. It uses inputs and execution results to compare behavioral differences among programs. There have been similar methods, such as the strict clustering approach in AlphaCode (Li et al., 2022), which samples a set of program inputs and then clustering programs with identical outputs into the same group. The final program is then selected from the largest cluster.
However, the idea of grouping programs by the ‘identicalness’ of outputs is not without fallacies, since programs rarely specialize in solving one single irreducible problem – they deal with a variety of inputs, conditions and mysterious cases. The result of this, where different solutions could have many common behaviors and some distinct behaviors, is referred to as the term ‘special-case similarity’ in the FrAngel paper (Shi et al., 2019). We consider a correct program solution which has multiple ‘special-case similar’ programs that are partially correct in different ways, for example:
  • One program behaving correctly on the general case (almost all) but missed a few edge cases
  • Another program got one edge case correct but didn’t manage to deal with the general case
  • Yet another program got all edge cases correct but crashed on the general case
  • A buggy program that behaves correctly on all available test cases but none of these tests trigger the bug (literally test coverage problem)
  • And many programs that turn out to be frenzy mixtures of all the above
If we had a pool of programs that contained the fully-correct program and an assortment of other programs that respected certain cases of the problem as aforementioned, it’d be obvious that the fully-correct would be decently ‘special-case similar’ to the rest of the programs, for their similar behavior on inputs. These execution outputs are programmatically obtained and automatically compared against each other without any human intervention or LLM calling required, the process of which sits at the core of our functional consensus algorithm.
Therefore, with functional consensus, where the solutions with common behaviors are promoted, we could intuitively expect the result to be a higher likelihood of a fully-correct program. Provided below is a hypothetical example demonstrating why functional consensus prevails:
Example
Consider the problem of finding all square roots in the complex domain of a non-zero real number (stored in float32). To get the answer right for all inputs, the function must consider 2 cases: A) non-zero numbers have 2 square roots; B) square roots of negative numbers are imaginary. 10 candidate functions are sampled as below:
  • 5 results ($a_{1},a_{2},a_{3},a_{4},a_{5}$) only considered case A and got just positive inputs right. For negative numbers, they literally gave up and crashed.
  • 3 results ($b_{6},b_{7},b_{8}$) remembered to consider case B, gave 2 imaginary results for negative numbers, but forgot to do the same for positive numbers, returning only 1 square root therein.
  • Only 2 results ($c_{9},c_{10}$) considered all cases and returned correct results for all inputs.
If we pick the program through ‘clustering’, the final result would be one of the 5 results ($a_{1},a_{2},\ldots,a_{5}$) that only considered case A, which is evidently not the correct solution. But with functional consensus, the final result is vastly different, since we consider the similarity between the functions based on their behavior on different inputs. Without loss of generality, suppose that there are 2 test inputs $4.0$, $-9.0$, one for each of the 2 cases. We calculate the similarity as follows:
  • Programs $a_{i}$ got only $\sqrt{4.0}=[2.0,-2.0]$ right so each program here are similar with programs ($a_{1},a_{2},a_{3},a_{4},a_{5},c_{9},c_{10}$), scoring 7 points.
  • Since $b_{j}$ only went well with $\sqrt{-9.0}=[3.0i,-3.0i]$, programs here only score 5 points for case B with the ones ($b_{6},b_{7},b_{8},c_{9},c_{10}$).
  • Each program in $c_{k}$ gets 7 points for $\sqrt{4.0}$ with ($a_{1},a_{2},a_{3},a_{4},a_{5},c_{9},c_{10}$), and gets 5 points for $\sqrt{-9.0}$ with ($b_{6},b_{7},b_{8},c_{9},c_{10}$). Totals to 12 points.
The final result apparently leaned towards $c_{k}$ as the correct solution, even if their outputs as a whole weren’t even half as much as $a_{i}$ is. Through this example, we illustrate that functional consensus has the potential to identify the correct samples even at their minority, outperforming other methods such as self-consistency or clustering.
A.9 Supplementary Results
Token Usage